Operating Systems and Systems Integration

A quick shell programming quiz

1. Write a program to create a local account for yourself if you do not already have a local account.

You should use useradd.

2. Extend your program from the previous question to take any username as a parameter and then test to see if there is a local account in that name. If not, it creates a local account for that user.

Nick Urbanik <nicku(at)nicku.org>  ver. 1.1